what is the slope of a line passing through the points (24,3)and (1,5)?
Answer by stanbon(75887)

what is the slope of a line passing through
the points (24,3)and (1,5)?
-------------------
slope = m = (3-5)/(24-1) = -2/23
==================================
Cheers,
Stan H.